standfast01:

Which expression is an equivalent expression of 12x + 10 + 4y? A. 2(6x + 2y + 6) B. 3(4x + y + 10) C. 2(6x + 4y + 8) D. 2(6x + 5 + 2y)

adevanath:

need help

standfast01:

Yep

umm:

You pretty much multiply the number outside the parentheses by everything that's inside, so like - 2 × 6x = 12x, 2 × 5 = 10, 2 × 2y = 4y. Final answer: 2(6x + 5 + 2y).

standfast01:

Thx

umm:

You're welcome. C:

standfast01:

K thx :)